# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

### 24 Jan 2024 v3

@@ I Introduction
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein
@@ II.1 The travelling waves
$L\frac{\partial I^{c}_{n}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
(And, of course, in the framework of the approximation the derivative in (1a) should be understood as partial derivative.)
From (1a) and (3) we can obtain closed wave equation for the voltage $V$
@@
However, further on we will not use Eq. (4); it will be more convenient to use Eqs. (1a) and (3).
@@
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (3) to get
$ULI^{c}=-\frac{\partial V}{\partial Z}-\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
Substituting (6) into (1a) we obtain ordinary differential equation
@@
Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
$E=\frac{dV}{dZ}$ (8)
and considering $V$ as the independent variable, we may after taking into account that
@@
and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
@@
A simple mechanical interpretation of Eq. (10) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of (10).
@@
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton.
The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (10) follows
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (12)
The inequality (12) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by Eq. (10), with the l.h.s. being put equal to zero. Hence the speed of the sound is given by the equation
$\frac{1}{u^{2}L}=\frac{dQ}{dV}.$ (13)
Now we understand that (12) means
@@
The conditions (15) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
In the vicinity of $V_{1,2}$ Eq. (10) has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (16)
Taking into account the boundary conditions (15) we obtain
$E\sim\left(V-V_{1,2}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (17)
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (18)
Eq. (10) can be specified to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (19)
For brevity we’ll call Eq. (18) and (19) the cubic case. Not that in this case
$\frac{1}{LC_{0}u^{2}(V)}=\alpha+2\beta V+3\gamma V^{2}.$ (20)
@@ II.2 The kinks
With four boundary conditions on $E(V)$ supplied by (15) and (17) for the kink problem of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two being connected with it. One such connection we could obtain at much earlier stage. In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{U^{2}L}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (21)
@@
$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2}.$ (22)
In the cubic case, calculating the integral in (21) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right).$ (23)
@@
$E^{2}(V)=\psi^{2}(V-V_{1})^{2}(V-V_{2})^{2},$ (24)
@@
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (26)
The solution is
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (27)
After finding $V(Z)$ we can from (2) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2UL}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (28)
@@ II.3 The solitons
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0\Rightarrow E(V_{3})=0.$ (29)
Hence
$E^{2}\sim\left(V-V_{3}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{3}.$ (30)
@@
With three boundary conditions on $E(V)$ supplied by (15) and (29) for the soliton problem of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
@@
$\frac{1}{LC_{0}U^{2}}=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right).$
We also can specify (32) (generalize (24)) to
@@
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (18). In this case, excluding $V_{4}$, we obtain that (34) and (II.3) are modified to
@@
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (36b)
$\frac{1}{LC_{0}U^{2}}$ $=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right).$ (36c)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (36c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation (20) with $\gamma=0$, in line with our previous general statement that the soliton is supersonic.
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (37)
The solution of (37) is
@@
Similar to what was done in the previous subsection, we can use (2) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (39)
@@ III The dissipative kinks and shocks
In this case (1c) is modified to
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (42)
hence Eq. (3) takes the form
$L\frac{\partial I^{c}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}-\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial Z^{2}}.$ (43)
(we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s. of (42) we presented as the continuous second order derivative).
From (1a) and (43) we can obtain Eq. (4) modified due to the presence of ohmic resistors
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (44)
For the travelling wave, combining (1a) and (43), we obtain after integration closed equation for $V(Z)$
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=U^{2}LQ-V+F,$ (45)
where $F$ is the constant of integration.
@@
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
@@
$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
@@
When both $\beta>0$ and $\gamma>0$, the condition (49) means just $|V_{1}|>|V_{2}|$.
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The stable equilibrium waves we called shocks, and the unstable equilibrium - kinks. There are several important distinctions between these two types of waves.
Unstable equilibrium at both at the point $V=V_{1}$ and $V=V_{2}$ means that differentiating the r.h.s. of (45) we recover equation (12). The kink, similar to what was in the lossless case, is supersonic both with respect to the region before the kink and the region behind the kink. Since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation.
Stable equilibrium at the point $V=V_{2}$ means
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{2}}-1<0.$ (50)
Inequalities (50) and (12) (the latter only for $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
To clarify the distinction between the shocks and the kinks let us bring together the formulas for the travelling wave speed and the sound speed for the cubic case
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right),$ (51a)
$\frac{1}{LC_{0}u^{2}(V_{1,2})}=\alpha+2\beta V_{1,2}+3\gamma V_{1,2}^{2}.$ (51b)
We realize that for $\beta>0$ and $\gamma>0$, for $V_{1}$ and $V_{2}$ having the same sign the travelling wave is always the shock. For $V_{1}$ and $V_{2}$ having opposite signs, we can have both the kink and the shock, depending upon the parameters.
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L\frac{dQ(V)}{dV}-1.$ (52)
The boundary conditions (11) lead to the boundary conditions (15) for $E(V)$, like it was in the lossless case. Equation (52) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=\text{const},$ (53)
which contains additional term in the l.h.s. with respect to (16), but we still obtain the condition (17). We can also repeat verbatim the part of Section II.2 leading to Eqs. (21) and (22).
In the cubic case Eq. (52) becomes
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (54)
We regain Eq. (23), and there exists kogan4 a simple polynomial solution of (54) in the form (24), which we repeat here
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (55)
where $\psi$ is a constant (positive for positive $V_{1}$). Substituting (55) into (54) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (56a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4UL}{\psi R}.$ (56b)
Note that Eqs. (25a), (35a) and (56a) are all identical, hence all the waves we consider in this paper can exist only if $\gamma>0$. Note also that by integrating (54) we can obtain Abel equation of the second kind, and (56b) is the condition of the solvability of the equation polyanin .
Equation (55) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (27)). And the the shocks are of the same shape also. However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (56b) instead of (25b)) and so does the kink (the shock) speed, in spite of the validity in all the cases of Eq. (23).
The boundary in the phase space between the kinks and the shocks can be easily obtained from (III) in the case when $\beta=0$. Equating the speeds we obtain that this boundary corresponds to
@@
We need to emphasise the different meaning of the condition (56b) for the kinks and the shocks. For the kinks, (56b) can be understood as the constraint on $V_{2}$ (for a given value of $V_{1}$ and ohmic resistance $R$). And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (56b) is just the condition for it being described by an elementary function. But, of course, the shock with any other value of $V_{2}$ (satisfying (48) also exists.
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1.$ (58)
The solution of (58) satisfying the boundary conditions (15) and (29) is
$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right],$ (59)
where $\Delta V\equiv V_{1}-V_{2}$ (we assumed for the sake of definiteness that $V_{1},V_{2}>0$). Substituting (59) into (58) we obtain
@@
$\frac{UL}{R}$ $=\frac{5\chi}{24}\left(\Delta V\right)^{1/2}.$ (60b)
The function $E(V)$ being found, we can find the function $V(Z)$ by solving equation
$\frac{dV}{dX}=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right].$ (61)
The solution is
$V(Z)=V_{2}+\frac{\Delta V}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (62)
According to what was said after Eq. (III), Eq. (62) describes the shock. And we have to repeat here what was said after Eq. (III). Equation (60b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $V_{2}$.

### 24 Feb 2024 v5

@@ I Introduction
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein ; vainchtein2 ; kamchatnov
@@ II.1 The travelling waves
From (1a) and (2) we can obtain closed wave equation for the voltage $V$
$L\frac{d^{2}Q_{n}}{dt^{2}}=V_{n+1}-2V_{n}+V_{n-1}.$ (3)
@@
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This from (4) we obtain an ordinary differential equation
$U^{2}L\frac{d^{2}Q(V)}{dZ^{2}}=\frac{d^{2}V}{dZ^{2}}+\frac{1}{12}\frac{d^{4}V}{dZ^{4}}.$ (6)
Performing an integration we get
@@
One way to solve Eq. (7) is to perform the second integration to obtain
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}=U^{2}LQ(V)-V+F,$ (8)
where $F$ is the constant of integration.
However, in this Section we’ll use another method of solution of Eq. (7). After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, we are motivated to introduce the new dependent variable
$E=\frac{dV}{dZ}.$ (9)
Considering $V$ as the independent variable, taking into account that
@@
and dividing both parts by the common multiplier $E$ in both parts, we may reduce the order of (7) and write it down as
@@
Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V$. These waves are described by Eq. (11), with the l.h.s. being put equal to zero. The speed of the sound is given by the equation
$\frac{1}{u^{2}(V)}=L\frac{dQ}{dV}.$ (12)
Thus Eq. (11) can be presented as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1.$ (13)
Note that in the framework of the quasi-continuum approximation $E$ is just the voltage on the inductor. Thus the physical meaning of Eq. (13) is simple. Its solution gives the voltage on the inductor as the function of the voltage on the capacitor.
A simple mechanical interpretation of Eq. (13) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of the equation.
@@
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton. The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (13) follows
@@
Also, from (10a) and (14) follows that in the vicinity of $V_{1,2}$
$E^{2}\sim\left(V-V_{1,2}\right)^{2}.$ (17)
The conditions (16) and (17) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
@@ II.2 The kinks
With four boundary conditions specifying both the position and the velocity of the particle at $V=V_{1}$ and $V=V_{2}$, of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two are connected with it. One such constraint we could obtain at much earlier stage. By integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{U^{2}}=L\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (18)
@@
$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2},$ (19)
where $P_{1}(V)$ is some polynomial function.
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right)$ (20)
(for brevity we’ll call Eq. (20) the cubic case), then
$\frac{1}{u^{2}(V)}$ $=LC_{0}\left(\alpha+2\beta V+3\gamma V^{2}\right),$ (21a)
$\frac{1}{U^{2}}$ $=LC_{0}\left[\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right)\right].$ (21b)
@@
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (22)
where $\psi$ is a constant (positive, if $V_{1}>V_{2}$, and negative in the opposite case). Substituting (22) into (13) and using Vieta’s formula we obtain after simple algebra
@@
Thus the necessary condition for the existence of the kink is the inequality $\gamma>0$, that is the speed of sound $u^{2}(V)$ should be a convex function of $V$.
To find $V(Z)$ we have to solve Eq. (9). The solution is
$V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (24)
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=-\frac{\left(V_{1}-V_{2}\right)}{2UL}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (25)
@@ II.3 The solitons
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0,\hskip 28.45274pt\left(\frac{d^{2}V}{dZ^{2}}\right)_{V=V_{3}}\neq 0.$ (26)
Hence from (9) and (10a) follows that in the vicinity of $V_{3}$
$E^{2}\sim\left(V-V_{3}\right).$ (27)
@@
With the position and the velocity of the particle at $V=V_{1}$ and its position at $V=V_{3}$ being specified, of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
@@
where $P_{2}(V)$ is some polynomial function.
@@
$\frac{1}{U^{2}}=LC_{0}\left[\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right)\right].$
We also can specify (29) to
@@
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (20). In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (31) and (II.3) are modified to
@@
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta.$ (33b)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
To find $V(Z)$ we have to solve Eq. (9). The solution is
@@
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{V_{3}-V_{1}}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (35)
@@ III The dissipative kinks and shocks
In this case the circuit equations are modified to
$L\frac{dI_{n,n+1}}{dt}$ $=\left(1+R_{C}C\frac{d}{dt}\right)\left(V_{n}-V_{n+1}\right),$ (38a)
$I^{c}_{n}$ $=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(I_{n-1,n}-I_{n,n+1}\right)$ (38b)
(Eq. (1a) remains as it was). Hence (3) is modified to
$L\frac{d^{2}Q_{n}}{dt^{2}}=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(1+R_{C}C\frac{d}{dt}\right)$
$\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (39)
and (4) is modified due to the presence of ohmic resistors to
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\tau_{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}},$ (40)
where
$\tau_{R}=\frac{L}{R_{L}}+R_{C}C.$ (41)
In deriving (40) we assumed that both the influence of the higher order derivatives with respect to $Z$ and the dissipation are small, so we ignored the terms which which correspond to the products of these effects.
For the travelling wave we obtain from (40) after double integration the generalization of (8)
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=U^{2}LQ(V)-V+F.$ (42)
@@
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (43)
@@
$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q(V)+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (44)
@@
For $\gamma>0,\beta\geq 0$, the condition (46) means just $|V_{1}|>|V_{2}|$.
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. (We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
Differentiating (44) we obtain the r.h.s. of (49) (see below)
$-\frac{d^{2}\Pi(V)}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1;$ (47)
the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background. Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{2}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
Let us return to (II.2) giving the sound and the kink/shock speeds for the cubic case. The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$. Consider the case $\gamma>0,\beta\geq 0$. If $V_{1}$ and $V_{2}$ have the same sign, the travelling wave is always the shock. In the opposite case, the wave can be either the kink or the shock, depending upon $V_{1}$ and $V_{2}$. The boundary in the phase space between the kinks and the shocks is especially simple for $\beta=0$
@@
The most important distinction of the dissipative kinks from the shocks: since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation (see below). The shocks don’t need such find tuning – they are generic.
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=\frac{U^{2}}{u^{2}(V)}-1.$ (49)
At this stage we can repeat verbatim the part of Section II.2 leading to (22), hence the dissipative kinks and the shocks are of the same shape (given by (24)) as the kinks in the lossless line.
However, Eq. (II.2) is generalized to
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (50a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (50b)
hence the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (50b) instead of (23b)) and so does the kink (the shock) speed and amplitude, in spite of the validity in all the cases of Eq. (21b).
We need to emphasise the different meaning of the condition (50b) for the kinks and the shocks. For the kinks, (50b) is the constraint on $V_{1},V_{2},R$. And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (50b) is just the condition for it being described by an elementary function. But, of course, the shock with any other values of $V_{1},V_{2},R$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (45)). Note also that by integrating (49) we can obtain Abel equation of the second kind, and (50b) is the condition of the solvability of the equation polyanin .
@@
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1$ (51)
(we assume for the sake of definiteness that $\beta>0$ and $V_{1},V_{2}>0$). The elementary solution of (51) satisfying the boundary conditions (16) is of the form kogan4
$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(V_{1}-V_{2}\right)^{1/2}\right].$ (52)
Substituting (52) into (51) we obtain
@@
$U\tau_{R}$ $=\frac{5\chi}{24}\left(V_{1}-V_{2}\right)^{1/2}.$ (53b)
Similar to Eq. (50b) in the case of shocks, Eq. (53b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $\Delta V$.
The function $E(V)$ being found, we can find the function $V(Z)$ by solving Eq. (9). The solution is
$V_{2}(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (54)